PROBLEM #1_________(10 PTS.) Assume that you are an agent for the Seattle Mariner’s next superstar and that you are negotiating with a number of major league baseball teams concerning the superstar’s next contract. The Mariners are hopeful that the superstar will sign with them and have offered a package containing the following:- Signing Bonus at the start of Year 1: $10,000,000 -7-Year contract paying $12,000,000 per year -A lump-sum payment made at the end of 10 years of $5,000,000

The Los Angeles Dodgers have countered with an attractive package of their own which includes:-Signing Bonus at the start of Year 1: $1,000,000 -7-Year contract paying $15,000,000 per yearREQUIRED:1. Which contract is the best “financial deal?

Explain. Show Computations. Use an interest rate of 10%.
Answer #1

Answer: The contract of The Los Angeles Dodgers is the best financial deal.

Calculations and explanations:

To determine the best financial deal we will find the present value of both the offers. The offer with the higher present value will be better.

Present value computation of offer 1 i.e. offer by The Mariners: (all numbers rounded to the nearest whole number i.e. 0 decimals):

Year (n) Amount 1+r PVIF = 1/(1+r)^n PV = Amount * PVIF
0    10,000,000           1.10                        1.0000                       10,000,000
                1    12,000,000                        0.9091                       10,909,091
                2    12,000,000                        0.8264                          9,917,355
                3    12,000,000                        0.7513                          9,015,778
                4    12,000,000                        0.6830                          8,196,161
                5    12,000,000                        0.6209                          7,451,056
                6    12,000,000                        0.5645                          6,773,687
                7    12,000,000                        0.5132                          6,157,897
                8                     -                          0.4665                                         -  
                9                     -                          0.4241                                         -  
              10      5,000,000                        0.3855                          1,927,716
Total PV                       70,348,742

PV of offer of The Los Angeles Dodgers :

Year (n) Amount 1+r PVIF = 1/(1+r)^n PV = Amount * PVIF
0      1,000,000           1.10                        1.0000                          1,000,000
                1    15,000,000                        0.9091                       13,636,364
                2    15,000,000                        0.8264                       12,396,694
                3    15,000,000                        0.7513                       11,269,722
                4    15,000,000                        0.6830                       10,245,202
                5    15,000,000                        0.6209                          9,313,820
                6    15,000,000                        0.5645                          8,467,109
                7    15,000,000                        0.5132                          7,697,372
Total PV                       74,026,282

We can see that present value of The Los Angeles Dodgers' offer is more and hence the contract of The Los Angeles Dodgers is the best financial deal.